# MatFormer: Nested Transformer for Elastic Inference

### v1

@@ E.2 Traditional Image Retrieval Evaluation
TriviaQA (EM) 4.63 3.87 4.87 4.55 6.11 5.63 8.09 6.48
NaturalQuestions (EM) 0.61 0.58 0.8 0.89 0.94 1.16 1.66 1.25
WebQuestions (EM) 2.31 1.62 2.26 2.02 2.85 2.31 2.85 2.56
LAMBADA 2.1 1.65 2.6 2.1 3.94 2.93 3.49 3.49
HellaSwag 32.12 31.57 32.83 32.16 33.8 33.48 36.21 35.08
StoryCloze 61.25 60.98 61.36 61.46 63.66 62.21 64.24 64.08
WSC 57.54 64.91 61.4 62.11 66.32 62.11 61.05 63.16
WinoGrande 52.33 51.38 49.09 50.99 52.64 50.36 53.12 52.64
Winograd 60.07 63.74 60.07 62.27 67.4 61.54 68.5 63.74
SQuAD v2 (EM) 21.7 21.85 25.8 19.71 24.69 21.85 23.08 18.28
RACE-H 29.85 29.45 29.47 29.79 30.56 29.79 30.7 30.02
RACE-M 37.53 37.6 37.33 38.93 40.39 39.62 40.95 39.21
PIQA 61.26 61.53 61.48 62.08 60.99 63.22 63.17 63.71
ARC-C 23.04 22.7 24.06 22.35 24.49 22.18 23.72 23.63
ARC-E 45.83 44.44 46.3 45.62 47.73 47.85 51.73 49.12
OpenBookQA 37.2 36.4 37 37.8 36.4 39.2 41 38.4
BoolQ 52.39 52.69 56.12 52.05 50.28 51.28 54.98 47.95
COPA 67 62 73 63 71 63 67 66
RTE 52.35 53.07 53.43 52.71 52.35 52.71 52.35 51.99
@@
MultiRC (F1) 45.63 46.02 54.4 46.38 52.79 49.28 52.34 41.71
ReCoRD 57.58 58.65 59.31 59.71 60.87 61 63.42 61.77
CB 42.86 42.86 44.64 42.86 44.64 42.86 42.86 42.86
ANLI-R1 32.6 32.5 31.7 33.1 31.4 32.3 32.5 32.6
ANLI-R2 30.7 30.7 28.4 30.5 30.4 30.6 31.2 31.8
ANLI-R3 30.83 30.67 30.08 30.75 30.83 30.67 30.92 30.75
Average 38.02 38.11 39.04 38.2 39.8 38.71 40.33 38.83
Avg over GEN Taks 6.27 5.91 7.27 5.85 7.71 6.78 7.84 6.41
Avg over RANK Tasks 45.59 45.77 46.61 45.9 47.44 46.31 48.06 46.55
Dev set log pplx 3.205 3.217 3.16 3.16 3.096 3.11 3.023 3.06
@@
TriviaQA (EM) 9.26 6.62 10.82 9.78 11.07 11.72 13.31 13.76
NaturalQuestions (EM) 1.66 0.89 1.69 1.58 2.24 2.38 2.66 2.74
WebQuestions (EM) 3.89 3.35 4.08 4.18 3.74 4.43 4.08 5.31
LAMBADA 3.2 8.25 6.97 10.83 8.19 10.44 14.03 10.83
HellaSwag 36.11 36.64 38.26 37.7 40.63 39.64 43.4 42.55
StoryCloze 64.78 65.26 66.33 66.17 68.25 67.13 71.25 69.64
WSC 66.32 65.96 63.16 64.21 69.82 69.12 70.53 68.42
WinoGrande 52.17 51.54 52.25 52.57 55.17 52.96 54.14 54.62
Winograd 68.13 69.23 67.03 71.43 71.06 70.33 72.16 72.89
SQuAD v2 (EM) 29.9 23.79 29.07 25.51 25.07 26.39 33.41 28.46
RACE-H 30.39 30.76 31.93 31.88 32.53 31.88 33.79 32.73
RACE-M 40.95 40.95 42.06 41.16 42.27 42.55 44.64 42.48
PIQA 64.04 63.98 64.64 64.91 65.45 65.23 67.25 66.21
ARC-C 24.49 24.15 26.71 24.91 26.71 26.54 27.13 27.47
ARC-E 52.15 51.01 53.66 52.95 56.27 54.92 57.11 56.57
OpenBookQA 38.2 40.4 40.8 41.2 42.8 40.8 43 42
BoolQ 52.63 50.31 51.9 47.8 56.73 50.15 55.6 48.41
COPA 68 73 68 73 71 73 73 76
RTE 51.62 51.99 52.71 52.35 51.62 51.99 53.07 52.71
WiC 47.34 47.18 47.34 47.18 47.34 47.18 47.34 47.18
MultiRC (F1) 44.37 51.32 52.11 50.46 54.7 53 37.58 47.16
ReCoRD 63.52 64.27 65.03 65.36 67.55 66.53 69.56 68.03
CB 42.86 37.5 42.86 42.86 42.86 42.86 46.43 39.29
ANLI-R1 30.9 31.8 33.7 32.1 31.7 32.2 32.6 32.4
ANLI-R2 31.8 31.5 31.5 30.9 31.1 30.6 30.4 30.8
ANLI-R3 32 30.25 32.83 30.17 30.75 30 30.58 30.25
Average 40.41 40.46 41.44 41.27 42.56 42.08 43.39 42.65
Avg over GEN Taks 9.58 8.58 10.53 10.38 10.06 11.07 13.5 12.22
Avg over RANK Tasks 47.75 48.05 48.8 48.63 50.3 49.46 50.5 49.9
Dev set log pplx 3.017 3.03 2.971 2.98 2.91 2.928 2.84 2.874
@@
Table 14: Downstream Eval numbers and development set log perplexity loss on 1.3B model size granularities.
TriviaQA (EM) 11.92 12 14.68 13.09 16.48 14.91 20.14 17.62
NaturalQuestions (EM) 1.88 2.19 2.24 2.47 3.07 2.99 4.79 4.13
WebQuestions (EM) 3.84 5.02 4.72 5.36 5.07 5.76 6.05 6.15
LAMBADA 7.3 9.94 13.55 12.34 17.97 13.51 22.65 19.21
HellaSwag 40.53 40.35 42.86 42.5 46 44.48 49.78 47.69
StoryCloze 67.29 68.2 69.75 69.91 72.37 71.14 73.81 72.8
WSC 64.56 65.96 64.91 69.12 67.72 69.82 72.63 69.82
WinoGrande 55.8 53.99 56.67 55.25 56.12 57.7 58.25 58.41
Winograd 71.06 68.5 67.77 70.7 73.99 70.33 72.53 72.89
SQuAD v2 (EM) 29.63 35.47 28.85 34.64 36.55 34.47 39.48 36.39
RACE-H 32.19 33.19 33.08 34.39 34.48 35.11 36.59 35.25
RACE-M 43.8 44.22 44.22 45.96 47.7 45.75 50.07 46.59
PIQA 66.49 64.36 66.05 66.38 67.52 66.97 69.1 67.68
ARC-C 27.99 25.77 27.65 27.22 29.01 28.75 30.55 31.48
ARC-E 56.44 54.08 58.54 57.03 59.85 58.84 63.26 61.83
OpenBookQA 41.4 42.2 41 42 43.4 42.8 44.8 45.4
BoolQ 52.57 49.85 54.86 52.42 53.76 56.06 55.35 53.52
COPA 70 75 69 77 74 74 77 75
RTE 52.35 53.07 53.07 52.35 54.15 53.43 52.35 49.82
WiC 47.34 47.34 47.18 47.34 47.34 47.34 48.43 47.02
MultiRC (F1) 42.98 46.69 43.82 49.09 45.29 48.2 40.99 46.42
ReCoRD 67.32 67 69.02 68.61 71.13 70.26 73.4 71.49
CB 42.86 44.64 46.43 42.86 48.21 44.64 42.86 37.5
ANLI-R1 32.5 33.5 31.9 33.8 33 33.3 32.4 32.1
ANLI-R2 30.3 34.7 30.5 34.6 30.6 33.1 31.5 33.5
ANLI-R3 30.5 33.17 31.5 33.67 31.33 33.5 32.58 33.67
Average 41.96 42.71 42.84 43.85 44.85 44.51 46.21 45.13
Avg over GEN Taks 10.91 12.92 12.81 13.58 15.83 14.33 18.62 16.7
Avg over RANK Tasks 49.35 49.8 49.99 51.06 51.76 51.69 52.77 51.9
Dev set log pplx 2.90 2.923 2.856 2.867 2.79 2.81 2.718 2.76
Table 15: Downstream Eval numbers and development set log perplexity loss on 2.6B model size granularities.
TriviaQA (EM) 18.58 18.64 19.83 21.41 25.17 24.9 28.84 28.01
NaturalQuestions (EM) 3.05 3.13 3.19 3.66 4.76 4.24 6.73 5.01
WebQuestions (EM) 5.61 6.74 4.43 6.3 6.1 6.74 8.27 7.78
LAMBADA 18.46 13.74 29.92 19.89 27.34 24.84 27.94 29.98
HellaSwag 46.41 46.01 49.04 48.94 52.87 52.2 57.14 55.33
StoryCloze 72.26 72.1 73.54 73.22 75.09 75.04 77.02 75.79
WSC 71.23 69.82 70.88 71.58 75.09 74.39 80 77.54
WinoGrande 56.83 57.85 57.62 56.91 60.93 59.19 62.19 59.59
Winograd 76.56 71.43 72.89 74.36 76.56 74.73 81.68 78.75
SQuAD v2 (EM) 34.89 37.97 34.33 40.07 34.89 42.24 43.47 42.59
RACE-H 33.62 34.76 35.59 35.85 36.91 36.82 38.91 37.28
RACE-M 47.63 47.49 49.44 49.51 50.77 50.07 53.34 51.67
PIQA 67.74 67.79 68.39 68.28 69.21 69.59 71.49 71.11
ARC-C 29.95 30.29 31.83 31.91 32.51 34.22 35.67 35.41
ARC-E 60.82 59.97 61.2 62.42 63.51 64.56 67.76 64.86
OpenBookQA 45.6 43.8 45.4 44.8 49 46.4 49 49.4
BoolQ 53.58 52.87 53.15 53.52 59.36 54.89 60.8 57.22
COPA 74 74 77 76 75 78 82 81
RTE 49.1 53.07 49.82 54.15 48.01 54.51 48.01 52.35
WiC 47.34 47.34 47.18 47.34 47.34 47.18 47.02 47.49
MultiRC (F1) 43.4 52.28 43.65 51.64 46.99 53.7 39.24 53.77
ReCoRD 71.34 71.9 72.79 72.97 74.86 74.57 76.71 75.32
CB 28.57 44.64 46.43 46.43 41.07 50 50 44.64
ANLI-R1 32.4 32.3 30.4 32.3 32.5 32.1 31.2 31.5
ANLI-R2 30.4 30.1 30.6 31 30.1 30.2 31.7 30.8
ANLI-R3 30.75 30.83 31.25 31 33.5 30.92 32 31.92
Average 44.23 45.03 45.76 46.36 47.29 47.93 49.54 49.08
Avg over GEN Taks 16.12 16.04 18.34 18.26 19.66 20.59 23.05 22.68
Avg over RANK Tasks 50.93 51.94 52.29 53.05 53.86 54.44 55.85 55.37
Dev set log pplx 2.77 2.787 2.722 2.732 2.66 2.68 2.592 2.63
Table 16: Downstream eval numbers and development set log perplexity on 2.6B MatLM Mix‘n’Match granularities. For original granularities, please refer to Table 15. First row represents the non-embedding parameters of the model.
Downstream Task 830M 1B 1.11B 1.32B 1.43B 1.55B 1.65B
TriviaQA (EM) 18.89 22.43 23.8 25.77 26.26 26.15 26.6
NaturalQuestions (EM) 3.49 3.77 4.02 4.07 4.46 4.65 5.12
WebQuestions (EM) 5.95 6.1 6.64 6.69 6.94 6.69 6.69
LAMBADA 16.34 20.16 23.07 24.8 24.32 25.87 29.13
HellaSwag 47.98 50.46 51.29 52.78 53.75 54.16 54.56
StoryCloze 73.01 73.33 74.83 75.2 75.68 75.41 75.63
WSC 70.88 70.53 74.04 72.98 74.74 73.33 77.19
WinoGrande 57.85 58.88 60.93 58.88 59.67 60.06 59.91
Winograd 73.26 73.26 76.19 74.36 76.56 77.66 78.02
SQuAD v2 (EM) 36.49 39.72 38.05 41.33 41.08 40.26 41.36
RACE-H 34.71 35.93 35.48 36.74 36.62 36.22 36.96
RACE-M 46.59 48.89 49.44 50.28 50.42 51.32 50.91
PIQA 68.5 69.04 69.53 70.4 70.46 70.51 70.29
ARC-C 31.06 33.11 33.19 34.81 35.75 35.84 34.56
ARC-E 62.29 62.58 62.63 64.86 65.99 65.49 64.69
OpenBookQA 44.6 46.2 46.8 47 47.4 47.4 47.6
BoolQ 54.86 55.08 54.46 55.78 58.38 57.19 56.88
COPA 76 76 75 80 77 80 80
RTE 53.43 53.79 53.79 52.71 53.79 54.51 53.79
WiC 47.34 47.34 47.18 47.34 47.18 47.34 48.12
MultiRC (F1) 53.34 53.85 52.97 54.23 57.57 55.09 54.91
ReCoRD 72.21 73.25 73.98 74.43 74.72 75.05 75.37
CB 48.21 46.43 48.21 50 50 44.64 55.36
ANLI-R1 32.4 32.1 32 32.4 32.3 31.4 32.4
ANLI-R2 30.5 30.6 30.6 30.6 30.7 30.4 31.4
ANLI-R3 31.17 31.17 31.17 31.5 31 31.5 31.33
Average 45.82 46.69 47.28 48.07 48.57 48.39 49.18
Avg over GEN Taks 16.23 18.44 19.12 20.53 20.61 20.72 21.78
Avg over RANK Tasks 52.87 53.42 53.99 54.63 55.22 54.98 55.71
Dev set log pplx 2.774 2.729 2.706 2.68 2.675 2.663 2.65

### v2

@@ G.2 Traditional Image Retrieval Evaluation
TriviaQA (EM) 4.63 5.67 4.87 6.78 6.11 7.51 8.09 7.76
NaturalQuestions (EM) 0.61 1.19 0.80 1.33 0.94 1.41 1.66 1.33
WebQuestions (EM) 2.31 2.61 2.26 2.56 2.85 2.71 2.85 3.20
LAMBADA 2.10 1.82 2.60 1.98 3.94 2.60 3.49 3.71
HellaSwag 32.12 32.69 32.83 34.05 33.80 34.88 36.21 36.36
StoryCloze 61.25 61.46 61.36 62.21 63.66 63.23 64.24 64.35
WSC 57.54 62.81 61.40 62.11 66.32 62.46 61.05 62.81
WinoGrande 52.33 51.46 49.09 51.78 52.64 50.20 53.12 51.85
Winograd 60.07 61.17 60.07 63.37 67.40 64.47 68.50 64.84
RACE-H 29.85 29.65 29.47 30.10 30.56 30.33 30.70 30.93
RACE-M 37.53 38.23 37.33 39.28 40.39 39.62 40.95 40.67
PIQA 61.26 61.43 61.48 62.40 60.99 62.79 63.17 63.44
ARC-C 23.04 24.23 24.06 24.32 24.49 25.43 23.72 24.91
ARC-E 45.83 46.42 46.30 47.35 47.73 49.54 51.73 50.72
OpenBookQA 37.20 37.40 37.00 38.60 36.40 38.80 41.00 39.20
BoolQ 52.39 51.77 56.12 51.96 50.28 53.52 54.98 55.17
COPA 67.00 65.00 73.00 64.00 71.00 63.00 67.00 70.00
RTE 52.35 55.23 53.43 50.54 52.35 49.46 52.35 52.35
@@
MultiRC 54.93 53.01 50.89 53.11 52.62 52.10 52.33 52.89
RECORD 57.58 59.93 59.31 61.06 60.87 62.16 63.42 63.51
CB 42.86 41.07 44.64 46.43 44.64 37.50 42.86 35.71
ANLI-RI 32.60 32.90 31.70 32.30 31.40 32.40 32.50 32.30
ANLI-R2 30.70 33.70 28.40 30.50 30.40 30.90 31.20 31.90
ANLI-R3 30.83 33.17 30.08 32.50 30.83 31.58 30.92 31.25
Average 39.05 39.64 39.43 39.91 40.40 39.83 41.01 40.83
Loss 3.205 3.162 3.16 3.107 3.096 3.06 3.023 3.02
@@
TriviaQA (EM) 6.62 10.50 9.78 11.61 11.72 13.15 13.31 15.15
NaturalQuestions (EM) 0.89 1.97 1.58 1.97 2.38 2.33 2.66 2.52
WebQuestions (EM) 3.35 3.69 4.18 2.95 4.43 3.94 4.08 4.82
LAMBADA 8.25 6.89 10.83 7.51 10.44 9.55 14.03 9.96
HellaSwag 36.64 37.42 37.70 39.21 39.64 41.48 43.40 43.43
StoryCloze 65.26 65.10 66.17 67.08 67.13 68.47 71.25 70.12
WSC 65.96 65.96 64.21 67.02 69.12 69.82 70.53 67.72
WinoGrande 51.54 52.57 52.57 54.46 52.96 54.54 54.14 54.85
Winograd 69.23 64.84 71.43 67.77 70.33 71.79 72.16 73.26
RACE-H 30.76 31.82 31.88 32.96 31.88 33.45 33.79 33.42
RACE-M 40.95 41.85 41.16 42.76 42.55 44.43 44.64 44.57
PIQA 63.98 64.04 64.91 65.02 65.23 66.81 67.25 67.19
ARC-C 24.15 26.02 24.91 27.65 26.54 28.84 27.13 30.38
ARC-E 51.01 52.69 52.95 55.13 54.92 56.27 57.11 59.01
OpenBookQA 40.40 41.60 41.20 42.20 40.80 42.20 43.00 43.40
BoolQ 50.31 52.48 47.80 54.80 50.15 54.59 55.60 54.46
COPA 73.00 74.00 73.00 73.00 73.00 75.00 73.00 78.00
RTE 51.99 54.87 52.35 54.15 51.99 54.15 53.07 53.43
WiC 47.18 47.34 47.18 47.34 47.18 47.18 47.34 47.18
MultiRC 51.79 55.01 53.09 54.81 53.36 55.61 56.08 56.19
RECORD 64.27 65.16 65.36 65.99 66.53 67.74 69.56 68.72
CB 37.50 41.07 42.86 44.64 42.86 48.21 46.43 50.00
ANLI-RI 31.80 32.70 32.10 31.80 32.20 31.20 32.60 31.30
ANLI-R2 31.50 30.20 30.90 30.00 30.60 30.40 30.40 30.90
ANLI-R3 30.25 30.92 30.17 31.33 30.00 32.08 30.58 31.92
Average 42.58 43.34 43.35 44.23 44.01 45.42 45.83 46.11
ANLI-R3 30.83 33.17 30.08 32.50 30.83 31.58 30.92 31.25
Average 41.14 42.03 42.01 42.92 42.72 44.13 44.53 44.87
Loss 3.017 2.987 2.971 2.934 2.91 2.886 2.84 2.843
@@
Table 14: Downstream Eval numbers and development set log perplexity loss on 850M model size granularities for DynaBERT and OFA
Downstream Task DynaBERT-S OFA-S DynaBERT-M OFA-M DynaBERT-L OFA-L DynaBERT-XL OFA-XL
TriviaQA (EM) 8.52 9.96 9.57 10.60 13.14 11.35 14.55 13.34
NaturalQuestions (EM) 1.39 1.52 1.83 2.24 2.55 2.22 2.96 2.60
WebQuestions (EM) 3.30 4.28 3.74 4.63 4.68 4.33 5.46 4.72
LAMBADA 8.85 4.72 11.84 5.74 14.88 5.98 14.65 8.31
HellaSwag 37.86 37.22 39.26 39.33 41.15 41.26 43.39 43.04
StoryCloze 65.42 64.40 66.54 66.49 68.36 68.57 69.05 69.75
WSC 64.91 68.07 68.77 65.96 71.58 68.77 72.63 68.42
WinoGrande 55.72 52.09 55.72 56.04 57.30 55.33 59.35 55.33
Winograd 69.23 68.86 69.60 68.50 72.53 70.33 75.46 72.53
RACE-H 31.16 32.70 32.30 32.96 32.62 32.99 33.19 33.96
RACE-M 41.43 42.41 43.04 43.04 45.54 43.73 46.52 43.73
PIQA 63.71 63.49 65.67 65.23 65.89 66.49 67.14 66.76
ARC-C 24.66 24.74 25.94 26.62 26.62 27.22 28.50 29.27
ARC-E 52.44 52.78 54.46 54.46 57.41 56.78 58.54 58.96
OpenBookQA 39.60 39.60 39.40 42.00 41.40 42.40 42.00 42.40
BoolQ 54.56 52.11 52.81 52.17 50.55 54.13 51.28 54.92
COPA 70.00 68.00 72.00 72.00 75.00 71.00 73.00 74.00
RTE 53.07 52.71 54.15 52.71 49.10 52.71 46.21 52.71
WiC 47.34 47.34 47.34 47.34 47.02 47.34 47.34 47.34
MultiRC 51.73 54.41 53.01 54.93 55.42 55.18 55.18 55.42
RECORD 65.05 64.60 66.65 66.61 68.33 67.96 68.91 68.68
CB 41.07 42.86 41.07 42.86 41.07 46.43 41.07 44.64
ANLI-RI 32.60 32.50 31.70 32.50 32.60 32.30 32.10 32.10
ANLI-R2 30.60 30.60 30.60 30.50 31.00 30.80 31.70 30.70
ANLI-R3 30.42 30.67 31.08 30.67 32.00 30.58 31.58 30.75
Average 41.78 41.71 42.72 42.64 43.9 43.44 44.47 44.17
Loss 2.993 3.01 2.942 2.935 2.895 2.89 2.854 2.863
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims in abstract are backed up by experimental results in Section 4.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We discuss this throughout the draft.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A] .
  12. Justification: No theoretical results.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Implementation details are provided in Appendix B. Moreover, code to reproduce Section 4.2 will be been open-sourced at link.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [No]
  20. Justification: Implementation details are provided in Appendix B. While code to reproduce Section 4.2 has been open-sourced, the language model has been trained on proprietary data.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Implementation details are provided in Appendix B, with citation of the models upon which these models are built.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: This has not been done due to the computational expense of training language models from scratch.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: This information is provided in Appendix B.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Broader impacts have been discussed in Appendix A.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: Broader impacts have been discussed in Appendix A.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: No datasets or models are released with this paper.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: The vision models and data used in this have been public for a long time. The language models inherit the appropriate licenses of the Lamda [58] paper while we train on proprietary data.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: Code to reproduce experiments will be released for camera ready.
  53. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  54. 14.
    Crowdsourcing and Research with Human Subjects
  55. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  56. Answer: [N/A]
  57. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  58. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  59. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  60. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  61. Answer: [N/A]
  62. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  63. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.